OpenStudy (anonymous):

Identify the 31st term of an arithmetic sequence where a1 = 26 and a22 = -226

OpenStudy (anonymous):

use \[a _{n}=a _{1}+\left( n-1 \right)d\] find d and then \[a _{31}\]

OpenStudy (anonymous):

so? a22=26+(22-1)d

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

so a22=26+(21)d

OpenStudy (anonymous):

then what? add 26+(21)?

OpenStudy (anonymous):

-226=26+21 d

OpenStudy (anonymous):

find d

OpenStudy (anonymous):

-273?

OpenStudy (anonymous):

21 d=-226-26=-252 d=-252/26

OpenStudy (anonymous):

-9.69....

OpenStudy (anonymous):

d=-126/13 find \[a _{31}\]

OpenStudy (anonymous):

I am so lost now

OpenStudy (anonymous):

it is simplification only.

OpenStudy (anonymous):

got the answer. -334 thanks

OpenStudy (anonymous):

\[a _{31}=26+(31-1)\times \frac{ -126 }{ 13 }\] there is something wrong either in your calculations or in statement.

OpenStudy (anonymous):

sorry ,i am leaving now.